Calculate the period of a simple pendulum if the length is 12 inches.


1
Expert's answer
2022-01-10T09:14:59-0500
"T=2\pi\sqrt{\dfrac{l}{g}},""T=2\pi\sqrt{\dfrac{12\ in\times\dfrac{0.0254\ m}{1\ in}}{9.8\ \dfrac{m}{s^2}}}=1.11\ s."